Use appropriate substitution to reduce the following equation to a variable separable and then solve the given IVP: "dy\/dx= (3x+2y)\/(3x+2y+2), y(-1)=-1"

1
Expert's answer
2022-02-21T16:43:53-0500

Let us solve the differential equation

"\frac{dy}{dx}=\frac{3x+2y}{3x+2y+2},\ \ y(−1)=−1."

Let us use the transformation "u=3x+2y." Then "\frac{du}{dx}=3+2\frac{dy}{dx}," and hence "\frac{dy}{dx}=\frac{1}2(\frac{du}{dx}-3)."

It follows that

"\frac{1}2(\frac{du}{dx}-3)=\frac{u}{u+2}," and hence "\frac{du}{dx}=\frac{2u}{u+2}+3=\frac{5u+6}{u+2}."

We get the equation "{dx}=\frac{(u+2)du}{5u+6}," and hence "\int{dx}=\int\frac{(u+2)du}{5u+6}=\frac{1}5\int\frac{(5u+10)du}{5u+6}
=\frac{1}5\int(1+\frac{4}{5u+6})du=\frac{u}5+\frac{4}{25}\ln|5u+6|+C."

Therefore,

"x=\frac{3x+2y}5+\frac{4}{25}\ln|5(3x+2y)+6|+C"

Since "y(-1)=-1," we conclude that "-1=-1+\frac{4}{25}\ln 19+C," and thus "C=-\frac{4}{25}\ln 19."

We conclude that the solution of "\frac{dy}{dx}=\frac{3x+2y}{3x+2y+2},\ \ y(−1)=−1" is

"x=\frac{3x+2y}5+\frac{4}{25}\ln|15x+10y+6|-\frac{4}{25}\ln 19."